# Calibrating an Imperfect Auxiliary Predictor for Unobserved No-Purchase Choice

### v1

@@ Abstract
Firms typically are unable to observe and record some consumer actions—whether they purchased from a competitor or decide not to purchase, or even whether they considered the firm’s product for purchase. This missing data on consideration and the choice of the outside-option makes identification and estimation of market size and choice parameters, even under simple models such as the Multinomial Logit (MNL), very challenging, and is one of the central problems in the estimation of choice models in an observed-data setting. A number of methods have been proposed recently to estimate the MNL model with unobserved no-purchases, often making use of auxiliary, related, market-share, or aggregated data. We study a setting in which such a black-box auxiliary predictor returns an estimate for the outside-option probability but may be biased or mis-calibrated for various reasons: it might have been trained on a different channel/market; or on market-share data from previous periods that have lost relevancy; or by an externally trained machine-learning system or a foundational AI model. In this paper we leverage a structural identity that links the outside-option log-odds to the utility of observed choices and develop calibration methods that convert such imperfect auxiliary predictions into statistically valid estimates of unobserved no-purchase probabilities under two scenarios: First, under an affine miscalibration model in logit space, we show that a simple regression identifies the outside-option utility parameters and yields consistent recovery of no-purchase probabilities (without the help of any new data). Second, assuming the auxiliary method approximately preserves the order we give a rank-based calibration approach and finite-sample error bounds that separate auxiliary-predictor quality from the utility learning error based on the items in the choice set. Numerical experiments illustrate the gains for no-purchase estimation and downstream assortment decisions, and we discuss extensions to aggregating multiple auxiliary predictors.
@@ 1 Introduction
Almost all these estimation ideas however rely on some external data or market-share signals, such as aggregated marginal competitor information or market-share data, or data from some other market with better observations on no-purchases, or managerial inputs and manual calibration. Such external aggregated data however usually has a time-lag—market-share data for instance is usually provided by information brokers quarterly or even yearly (for instance market-share data at an airport by origin-destination); The competitor aggregated sales data used in (Li et al., 2025a) is supplied with a month lag. Other sources of data has similar relevancy concerns: recent advances in machine learning and generative AI have made it possible to obtain predictions for unobserved outcomes (e.g., no-purchase events or competitors’ market shares) where data from a large number of external sources is used (Gallegos et al., 2024; Chen et al., 2025; Park et al., 2023). As a result, an estimator trained or calibrated on lagged aggregates or different data sets can create a bias in estimations of the outside-option probability even if they preserve useful structure (e.g., correlation or approximate ranking agreement with the truth). In this paper we assume the availability of such an auxiliary predictor for the outside option, which may be informative yet miscalibrated in the focal environment (the context of the current period where decisions are made) and give methods for correcting certain types of bias that may be present in them.
When does our calibration approach work?
Our approach is based on a structural identity in choice models that links the outside-option log-odds to the utility of observed choices. This identity implies that if the auxiliary predictor is miscalibrated by a shift-and-scale transformation in logit space across assortments and covariates, then a simple regression on the observed information identifies the unobserved outside-utility coefficients and yields consistent recovery of outside-option probabilities. We then extend our approach beyond affine miscalibration to nearly monotone predictors: when the auxiliary predictor preserves the ordering of outside-option logits for most samples, we can estimate the outside-option utility parameters via a maximum rank correlation (MRC) criterion (Han, 1987). We provide finite-sample error bounds for both situations, separating (i) the quality of the auxiliary predictor from (ii) the error from learning inside utilities using purchase-only data. The box below further clarifies the conditions required by our approach.
@@
  • Learnable in-set utilities from purchase-only data: there must be within-assortment variation (and implicitly, assortments of size greater than one) so that relative utilities of the products in the assortment can be learned from the conditional likelihood.
@@ 2.1 Synthetic Data for Operations Management
Using synthetic/generated data is becoming common in operations management to address data scarcity, costly labeling, and constraints on data sharing, while still enabling rigorous model development, benchmarking, and counterfactual experimentation. In manufacturing, Buggineni et al. (2024) provide a comprehensive review and propose a four-stage “collect, preprocess, generate, and evaluate” framework. They show that carefully constructed observations can improve downstream tasks, and they emphasize the importance of exploring how synthetic and real data can be combined in future works. Examples of applications in operations include (i) simulation-driven time series and event logs that instantiate production systems to create public and reusable datasets (e.g., Chan et al., 2022; Pendyala et al., 2024), and (ii) digital twin generators that parameterize factories and emit synthetic telemetry for bottleneck analysis and policy testing at scale (e.g., Lopes et al., 2024).
Synthetic data is also used in supply chain management and retail analytics. Long et al. (2025) to synthesize various scenarios of demand, risk, inventory, and perform forecasting and network optimization. Empirically, time series generators can help demand prediction when real observations are scarce or imbalanced. For instance, Chatterjee et al. (2025) show that sequence synthesis based on GANs strengthens electric vehicle related demand forecasting. Methodologically, a line of research studies how to build general-purpose generators for tabular and sequential data such as Conditional Tabular GANs (CTGAN) and related models that widely used (Xu et al., 2019; Patki et al., 2016); For sequential data, time-series GANs (TimeGAN) and their variants have been show to perform well empirically (Yoon et al., 2019; Esteban et al., 2017).
@@ 2.2 Unobservable Choice Prediction and Calibration
  • Two-step and moment-based estimators to exploit identifiability structure in the multinomial-logit model with assortment sizes strictly greater than 1. Specifically, first estimate relative utilities from within-purchase shares and then calibrate the outside-option per market size, e.g., using risk-ratio or related moment conditions (Talluri, 2009); or decompose the likelihood into marginal and conditional components to consistently estimate with a completely censored alternative using single-firm sales data (Newman et al., 2014). In addition, (Li et al., 2025a) propose GMM procedures when the observed purchase data have been shaped by a firm’s own revenue-management controls (assortment or pricing decisions), thereby addressing optimization-induced endogeneity; they do not require external market-share data.
@@ 3 Problem Setup
Let $\mathcal{I}$ denote a finite set of items. At each interaction we observe an offered subset (assortment) $\mathcal{S}\subseteq\mathcal{I}$ and an observed context $X\neq\emptyset$, and a customer (randomly) chooses one option $I\in\mathcal{S}\cup\{0\}$. Here 0 indexes an outside-option, capturing no-purchase or purchase from a competitor.
@@
In this paper, we only observe transactions in which an inside item is purchased ($I\in\mathcal{S}$); equivalently, the observed samples are drawn from the conditional distribution of $(X,\mathcal{S},I)$ given $I\neq 0$. In addition, we can query an auxiliary predictor on the same inputs $(X,\mathcal{S})$ to obtain a prediction of the (unobsevable) outside-option probability $\tilde{p}_{0}(X,\mathcal{S})$. We treat this predictor as a black box: we do not require access to (or knowledge of) its training data or proprietary feature engineering, only that we can evaluate its output on the observed $(X_{k},\mathcal{S}_{k})$ pairs.
@@
for an unknown coefficient vector $\gamma^{*}\in\mathbb{R}^{d}$. In (Tomlinson and Benson, 2021) they take $z(X)=X$ (where $X$ can be market features or a competitor-context summary). More generally, $z(X)$ may collect embeddings from kernel basis functions (Yang et al., 2025) or from a neural network choice model trained on another dataset (e.g., the representation before the output layer (Zhang et al., 2025; Li et al., 2025b; Wang et al., 2023a, b)).
@@
The predictor may be provided by a generative AI system, a consulting firm, or another external source; however, it can be biased or inaccurate because it is not trained on our data or tailored to our model. A natural source of a predictor is an auxiliary dataset or system that records the full funnel (arrivals/sessions) so that non-purchases are observable. Examples include (i) a limited time window in which a platform collected detailed browsing/search logs before reverting to transaction-only storage, (ii) a partner channel (or a different market) where session outcomes, including “no purchase,” are logged, or (iii) third-party tools (consultants, survey-based choice predictors, or foundation-model-based systems) trained on external data. A predictor trained in such environments can then be queried on the $(X,\mathcal{S})$ pairs observed in the focal transactions, but it may be systematically biased due to domain mismatch or covariate shift.
@@ 4 Key Structural Insight for Bias Correction
This section highlights a structural identity implied by the choice model (1). It shows that the (unobserved) outside-option probability can be summarized by a single index that decomposes into (i) an $X$-only term coming from the outside utility and (ii) an “inclusive value” term coming from the attractiveness of the offered inside options. This identity is the key reason we can debias predictor outputs without ever observing no-purchase events.
@@ 5 Linearly Biased Predictor
Our assumption in the section is the following:
@@ 5.1 Linear Regression Calibration
$\hat{\gamma}\;=\;\frac{\hat{\theta}_{z}}{-\hat{\theta}_{s}},$
so we recover $\gamma$ without knowing $a^{*}$ or $b^{*}$. Mapping the recovered index $\hat{\eta}(X,\mathcal{S})=\hat{\gamma}^{\top}z(X)-s(X,\mathcal{S})$ through the logistic function then yields a calibrated $\hat{p}_{0}(X,\mathcal{S})$ aligned with the model structure.
@@
If the utilities $\{u_{i}(X)\}_{i\in\mathcal{I}}$ are known, we regress $y$ on $\big[1,\,z(X),\,s(X,\mathcal{S})\big]$ to estimate $(a,\theta_{z},\theta_{s})$ and set $\hat{\gamma}=\hat{\theta}_{z}/(-\hat{\theta}_{s})$. In practice, we only have estimated utilities $\{\hat{u}_{i}(X)\}_{i\in\mathcal{I}}$ learned from historical data, so we replace
@@ 6 Nearly Monotone Biased Predictor
as the synthetic logit and the true logit of the unobservable outside choice, respectively. We then assume
@@ 6.2 Analysis
$\big|\hat{p}_{0}(X,\mathcal{S})-p_{0}(X,\mathcal{S})\big|\;\leq\;\frac{1}{4}\Bigg[\|z(X)\|_{2}\Bigg(\frac{2C_{U}\sqrt{\dfrac{d+d\log(n/d)+\log(2/\delta)}{n}}+2L_{\mathrm{ac}}(2\tau_{s})^{\alpha_{\mathrm{up}}}+\delta_{0}}{(1-2\rho_{0})\,\kappa}\Bigg)^{\!1/\alpha_{\mathrm{down}}}\\[6.0pt] \qquad\qquad\qquad\qquad\qquad\qquad+\;\big|\hat{s}(X,\mathcal{S})-s(X,\mathcal{S})\big|\Bigg].$ (12)
@@
so improvements in utility learning ($\tau_{s}\downarrow$) and cleaner simulation ($\delta_{0}\downarrow$, $\rho_{0}\downarrow$) reduce this asymptotic floor.
@@ 6.3 Example: Perturbative Monotone Link
Hence, if Assumption 6 holds, Assumption 5 holds with $\rho_{0}=0$ and
@@
$\big|\hat{p}_{0}(X,\mathcal{S})-p_{0}(X,\mathcal{S})\big|\;\leq\;\frac{1}{4}\Bigg[\|z(X)\|_{2}\,\Bigg(\frac{2C_{U}\sqrt{\frac{d+d\log(n/d)+\log(2/\delta)}{n}}\;+\;2L_{\mathrm{ac}}\,(2\tau_{s})^{\alpha_{\mathrm{up}}}\;+\;L_{\mathrm{ac}}\,t_{\Delta}^{\alpha_{\mathrm{up}}}}{\kappa}\Bigg)^{\!\frac{1}{\alpha_{\mathrm{down}}}}\\[6.00006pt] +\;\big|\hat{s}(X,\mathcal{S})-s(X,\mathcal{S})\big|\Bigg].$ (14)
@@ 7.1 Assortment Optimization
and $C^{\mathrm{pert}}_{n}(\delta)$ is the calibration coefficient guaranteed by Corollary 2, i.e. the smallest quantity for which, uniformly over $\mathcal{S}$,
@@ 8.2.1 Data splitting and covariate shift
Table 1 summarizes the two splits. While the no-purchase frequency is similar across periods, covariates shift substantially (e.g., the average listed price increases from \$225.95 to \$279.48), making the historical predictor becomes systematically miscalibrated on the current period, creating a realistic need for structural calibration. Intuitively, on average, the same input price leads to a lower no-purchase probability in the Historical Set (which trains the predictor) than the Current Set (which provides the real data without no-purchase).
@@ C.4 Refined Bound for Theorem 2
$\|\hat{\theta}-\theta^{*}\|_{2}\leq\underbrace{\frac{2B_{w}}{\lambda_{0}}\sigma\sqrt{\frac{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}{n}}}_{\text{oracle OLS term}}\\ +\underbrace{\frac{(4B_{w}\sqrt{\tau}+\tau)(\sqrt{\tau}+2B_{w})+\lambda_{0}\sqrt{\tau}}{\lambda_{0}(\lambda_{0}-4B_{w}\sqrt{\tau}-\tau)\sqrt{n}}\,\|\tilde{y}\|_{2}}_{\text{perturbation due to }\hat{s}_{k}}.$ (41)
@@
$\|\tilde{y}\|_{2}^{2}\leq 4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}\\ +\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr)$ (42)
@@
$\|\tilde{y}\|_{2}^{2}\leq 4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}\\ +\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr).$ (42)
@@
$\|\tilde{y}\|_{2}^{2}\leq 4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}\\ +\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr),$ (44)

### v2

@@ Abstract
Firms typically are unable to observe and record some consumer actions—whether they purchase from a competitor or decide not to purchase, or even whether they considered the firm’s product for purchase. This missing data on consideration and the choice of the outside-option makes identification and estimation of market size and choice parameters, even under simple models such as the Multinomial Logit (MNL), very challenging, and is one of the central problems in the estimation of choice models in an observed-data setting. A number of methods have been proposed recently to estimate the choice model with unobserved no-purchases, often making use of auxiliary, related, market-share, or aggregated data. We study a setting in which such a black-box auxiliary predictor returns an estimate for the outside-option probability but may be biased or miscalibrated for various reasons: it might have been trained on a different channel/market; or on market-share data from previous periods that have lost relevance; or by an externally trained machine-learning system or a foundational AI model. In this paper we leverage a structural identity that links the outside-option log-odds to the utility of observed choices and develop calibration methods that convert such imperfect auxiliary predictions into statistically valid estimates of unobserved no-purchase probabilities under two scenarios: First, under an affine miscalibration model in logit space, we show that a simple regression identifies the outside-option utility parameters and yields consistent recovery of no-purchase probabilities (without the help of any new data). Second, assuming the auxiliary method approximately preserves the order, we give a rank-based calibration approach and finite-sample error bounds that separate auxiliary-predictor quality from the utility learning error based on the items in the choice set. Numerical experiments illustrate the gains for no-purchase estimation and downstream assortment decisions, and we discuss extensions to aggregating multiple auxiliary predictors.
@@ 1 Introduction
Almost all these estimation ideas, however, rely on some external data or market-share signals, such as aggregated marginal competitor information or market-share data, or data from some other market with better observations on no-purchases, or managerial inputs and manual calibration. Such external aggregated data, however, usually has a time-lag—market-share data for instance is usually provided by information brokers quarterly or even yearly (for instance market-share data at an airport by origin-destination); the competitor aggregated sales data used in Li et al. (2025a) is supplied with a month lag. Other sources of data have similar relevancy concerns: recent advances in machine learning and generative AI have made it possible to obtain predictions for unobserved outcomes (e.g., no-purchase events or competitors’ market shares) where data from a large number of external sources is used (Gallegos et al., 2024; Chen et al., 2025; Park et al., 2023). As a result, an estimator trained or calibrated on lagged aggregates or different data sets can create a bias in estimations of the outside-option probability even if they preserve useful structure (e.g., correlation or approximate ranking agreement with the truth). In this paper we assume the availability of such an auxiliary predictor for the outside option, which may be informative yet miscalibrated in the focal environment (the context of the current period where decisions are made) and give methods for correcting certain types of bias that may be present in them.
Our proposed approach is based on a structural identity in choice models that links the outside-option log-odds to the utility of observed choices. This identity implies that if the auxiliary predictor is miscalibrated by a shift-and-scale transformation in logit space across assortments and covariates, then a simple regression on the observed information identifies the unobserved outside-utility coefficients and yields consistent recovery of outside-option probabilities. We then extend our approach beyond affine miscalibration to nearly monotone predictors: when the auxiliary predictor preserves the ordering of outside-option logits for most samples, we can estimate the outside-option utility parameters via a maximum rank correlation (MRC) criterion (Han, 1987). We provide finite-sample error bounds for both situations, separating (i) the quality of the auxiliary predictor from (ii) the error from learning inside utilities of observed choices using purchase-only data.
@@
  • Learnable utilities from purchase-only data: there must be within-assortment variation (and implicitly, assortments of size greater than one) so that utilities of the observable choices in the assortment can be learned from the conditional likelihood.
@@ 2.1 Synthetic Data for Operations Management
Using synthetic/predicted data is becoming common in operations management to address data scarcity, costly labeling, and constraints on data sharing, while still enabling rigorous model development, benchmarking, and counterfactual experimentation. In manufacturing, Buggineni et al. (2024) provide a comprehensive review and propose a four-stage “collect, preprocess, generate, and evaluate” framework. They show that carefully constructed observations can improve downstream tasks, and they emphasize the importance of exploring how synthetic and real data can be combined in future works. Examples of applications in operations include (i) simulation-driven time series and event logs that instantiate production systems to create public and reusable datasets (e.g., Chan et al., 2022; Pendyala et al., 2024), and (ii) digital twin generators that parameterize factories and emit synthetic telemetry for bottleneck analysis and policy testing at scale (e.g., Lopes et al., 2024).
Synthetic data is also used in supply chain management and retail analytics. Long et al. (2025) synthesize various scenarios of demand, risk, inventory, and perform forecasting and network optimization. Empirically, time series generators can help demand prediction when real observations are scarce or imbalanced. For instance, Chatterjee et al. (2025) show that sequence synthesis based on GANs strengthens electric vehicle related demand forecasting. Methodologically, a line of research studies how to build general-purpose generators for tabular and sequential data such as Conditional Tabular GANs (CTGAN) and related models that are widely used (Xu et al., 2019; Patki et al., 2016); For sequential data, time-series GANs (TimeGAN) and their variants have been shown to perform well empirically (Yoon et al., 2019; Esteban et al., 2017).
@@ 2.2 Unobservable Choice Prediction and Calibration
  • Two-step and moment-based estimators to exploit identifiability structure in the multinomial-logit model with assortment sizes strictly greater than 1. Specifically, first estimate relative utilities from within-purchase shares and then calibrate the outside-option per market size, e.g., using risk-ratio or related moment conditions (Talluri, 2009); or decompose the likelihood into marginal and conditional components to consistently estimate with a completely censored alternative using single-firm sales data (Newman et al., 2014). Relatedly, cho_ferguson_im_pekgun_2024_pom develop a transaction-level two-step estimator that avoids time-window aggregation (accommodating frequent price/attribute changes) by first estimating normalized utilities from the conditional (purchase-only) likelihood and then identifying lost sales/outside-option parameters using either external market-share information or mild utility restrictions. They further propose a robust procedure that uses a range of plausible market shares. In addition, Li et al. (2025a) propose GMM procedures when the observed purchase data have been shaped by a firm’s own revenue-management controls (assortment or pricing decisions), thereby addressing optimization-induced endogeneity; they do not require external market-share data.
@@ 3 Problem Setup
Let $\mathcal{I}$ denote a finite set of items. At each interaction we observe an offered subset (assortment) $\mathcal{S}\subseteq\mathcal{I}$ and an observed context $X$, and a customer (randomly) chooses one option $I\in\mathcal{S}\cup\{0\}$. Here 0 indexes an outside-option, capturing no-purchase or purchase from a competitor.
@@
In this paper, we only observe transactions in which an inside item is purchased ($I\in\mathcal{S}$); equivalently, the observed samples are drawn from the conditional distribution of $(X,\mathcal{S},I)$ given $I\neq 0$. In addition, we can query an auxiliary predictor on the same inputs $(X,\mathcal{S})$ to obtain a prediction of the (unobservable) outside-option probability $\tilde{p}_{0}(X,\mathcal{S})$. We treat this predictor as a black box: we do not require access to (or knowledge of) its training data or proprietary feature engineering, only that we can evaluate its output on the observed $(X_{k},\mathcal{S}_{k})$ pairs.
@@
for an unknown coefficient vector $\gamma^{*}\in\mathbb{R}^{d}$. In Tomlinson and Benson (2021) they take $z(X)=X$ (where $X$ can be market features or a competitor-context summary). More generally, $z(X)$ may collect embeddings from kernel basis functions (Yang et al., 2025) or from a neural network choice model trained on another dataset (e.g., the representation before the output layer (Zhang et al., 2025; Li et al., 2025b; Wang et al., 2023a, b)).
@@
The predictor may be provided by a generative AI system, a consulting firm, or another external source; however, it can be biased or inaccurate because it is not trained on our data or tailored to our model. A natural source of a predictor is an auxiliary dataset or system that records the full funnel (arrivals/sessions) so that non-purchases are observable. Examples include (i) a limited time window in which a platform collected detailed browsing/search logs before reverting to transaction-only storage, (ii) a partner channel (or a different market) where session outcomes, including “no-purchase,” are logged, or (iii) third-party tools (consultants, survey-based choice predictors, or foundation-model-based systems) trained on external data. A predictor trained in such environments can then be queried on the $(X,\mathcal{S})$ pairs observed in the focal transactions, but it may be systematically biased due to domain mismatch or covariate shift.
@@ 4 Key Structural Insight for Bias Correction
This section highlights a structural identity implied by the choice model (1). It shows that the (unobserved) outside-option probability can be summarized by a single index that decomposes into (i) an $X$-only term coming from the outside utility and (ii) an “inclusive value” term coming from the attractiveness of the inside options in the offered assortment. This identity is the key reason we can debias predictor outputs without ever observing no-purchase events.
@@ 5 Linearly Biased Predictor
Let $y(X,\mathcal{S})=\mathrm{logit}(\tilde{p}_{0}(X,\mathcal{S}))$ be the predictor’s logit. Our assumption in the section is the following:
@@ 5.1 Linear Regression Calibration
$\gamma^{*}\;=\;\frac{\theta_{z}}{-\theta_{s}},$
so we recover $\gamma^{*}$ without knowing $a^{*}$ or $b^{*}$. Based on the estimation $\hat{\gamma}$, mapping the recovered index $\hat{\eta}(X,\mathcal{S})=\hat{\gamma}^{\top}z(X)-s(X,\mathcal{S})$ through the logistic function then yields a calibrated $\hat{p}_{0}(X,\mathcal{S})$ aligned with the model structure.
@@
If the utilities $\{u_{i}(X)\}_{i\in\mathcal{I}}$ are known, we regress $y$ on $\big[1,\,z(X),\,s(X,\mathcal{S})\big]$ to estimate $(a,\theta_{z},\theta_{s})$ and set $\hat{\gamma}=\hat{\theta}_{z}/(-\hat{\theta}_{s})$ based on our estimations $(\hat{\theta}_{z},\hat{\theta}_{s})$. In practice, we only have estimated utilities $\{\hat{u}_{i}(X)\}_{i\in\mathcal{I}}$ learned from historical data, so we replace
@@ 6 Nearly Monotone Biased Predictor
as the predicted logit and the true logit of the unobservable outside choice, respectively. We then assume
@@ 6.2 Analysis
$\big|\hat{p}_{0}(X,\mathcal{S})-p_{0}(X,\mathcal{S})\big|\;\leq$ $\frac{1}{4}\Bigg[\|z(X)\|_{2}\Bigg(\frac{2C_{U}\sqrt{\dfrac{d+d\log(n/d)+\log(2/\delta)}{n}}+2L_{\mathrm{ac}}(2\tau_{s})^{\alpha_{\mathrm{up}}}+\delta_{0}}{(1-2\rho_{0})\,\kappa}\Bigg)^{\!1/\alpha_{\mathrm{down}}}$ (12)
$\qquad+\;\big|\hat{s}(X,\mathcal{S})-s(X,\mathcal{S})\big|\Bigg].$
@@
so improvements in utility learning ($\tau_{s}\downarrow$) and cleaner prediction ($\delta_{0}\downarrow$, $\rho_{0}\downarrow$) reduce this asymptotic floor.
@@ 6.3 Example: Perturbative Monotone Link
Hence, if Assumption 6 and  7 hold, Assumption 5 holds with $\rho_{0}=0$ and
@@
$\big|\hat{p}_{0}(X,\mathcal{S})-p_{0}(X,\mathcal{S})\big|\;\leq$ $\frac{1}{4}\Bigg[\|z(X)\|_{2}\,\Bigg(\frac{2C_{U}\sqrt{\frac{d+d\log(n/d)+\log(2/\delta)}{n}}\;+\;2L_{\mathrm{ac}}\,(2\tau_{s})^{\alpha_{\mathrm{up}}}\;+\;L_{\mathrm{ac}}\,t_{\Delta}^{\alpha_{\mathrm{up}}}}{\kappa}\Bigg)^{\!\frac{1}{\alpha_{\mathrm{down}}}}$ (14)
$\qquad+\;\big|\hat{s}(X,\mathcal{S})-s(X,\mathcal{S})\big|\Bigg].$
@@ 7.1 Assortment Optimization
and $C^{\mathrm{pert}}_{n}(\delta)$ is the calibration coefficient guaranteed by Corollary 2, i.e., the smallest quantity for which, uniformly over $\mathcal{S}$,
@@ 8.2.1 Data splitting and covariate shift
Table 1 summarizes the two splits. While the no-purchase frequency is similar across periods, covariates shift substantially (e.g., the average listed price increases from \$225.95 to \$279.48), making the historical predictor become systematically miscalibrated on the current period, creating a realistic need for structural calibration. Intuitively, on average, the same input price leads to a lower no-purchase probability in the Historical Set (which trains the predictor) than the Current Set (which provides the real data without no-purchase).
@@ C.4 Refined Bound for Theorem 2
$\|\hat{\theta}-\theta^{*}\|_{2}\leq$ $\underbrace{\frac{2B_{w}}{\lambda_{0}}\sigma\sqrt{\frac{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}{n}}}_{\text{oracle OLS term}}$ (41)
$+\underbrace{\frac{(4B_{w}\sqrt{\tau}+\tau)(\sqrt{\tau}+2B_{w})+\lambda_{0}\sqrt{\tau}}{\lambda_{0}(\lambda_{0}-4B_{w}\sqrt{\tau}-\tau)\sqrt{n}}\,\|\tilde{y}\|_{2}}_{\text{perturbation due to }\hat{s}_{k}}.$
@@
$\|\tilde{y}\|_{2}^{2}\leq$ $4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}$ (42)
$+\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr).$
@@
$\|\tilde{y}\|_{2}^{2}\leq$ $4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}$ (42)
$+\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr).$
@@
$\|\tilde{y}\|_{2}^{2}\leq$ $4nB_{w}^{2}\|\theta^{*}\|_{2}^{2}+\frac{4\sqrt{n}B_{w}\|\theta^{*}\|_{2}}{\lambda_{0}}\sigma\sqrt{1+C\log(4/\delta)+C\sqrt{\log(4/\delta)}}$ (44)
$+\sigma^{2}\bigl(n+C^{\prime}\sqrt{n\log(4/\delta)}+C^{\prime}\log(4/\delta)\bigr).$